OpenStudy (anonymous):

How do you fine the value of x in a kite when the only information your given is that one of the right triangles in the kite measures (7x-1)degrees?

OpenStudy (anonymous):

=360

OpenStudy (anonymous):

use the symmetric properties of a kite...

OpenStudy (anonymous):

4(7x-1)

OpenStudy (anonymous):

line symmetry...

OpenStudy (anonymous):

so 4(7x-1)=360 is the equation?

OpenStudy (anonymous):

do you have a picture that go along with what your question is?

OpenStudy (anonymous):

yeah, but i cant copy and paste onto here. its from an online summer school website im using.

OpenStudy (anonymous):

 so all 4 angles will total to 360 degrees. thus, 4(7x-1) = 360 28x - 4 = 360 x = 364/28 x = 13